# Online Learning of Whittle Indices for Restless Bandits with Non-Stationary Transition Kernels

### v1

@@ 4 Sliding Window-based Online Whittle Policy
Recall from earlier discussion (7) that to compute the Whittle index, we need to know transition kernels. In practice, transition kernels $P_{n,t}$ are unknown and non-stationary. In this section, we present Algorithm 1, an online approach for RMABs which adapts to unknown and non-stationary transition kernels.
input : State Space $\mathcal{S}$, Action Space $\mathcal{A}$, Reward Function $r_{n}(s,a)$, and the other domain knowledge $\epsilon_{n}$, $Z_{1}^{(n)},Z_{2}^{(n)},$ and $S_{0}^{(n)}(s,a)$ for all $(s,a)$ and arms $n$
1 DM decides window size $w_{n}$ for all $n\in[N]$, $\eta_{1}$, $\eta_{2}$ and initialize a Lagrange cost $\lambda^{(1)}$
@@
3      All arms starts with state $s_{n,0}$ for all $n\in[N]$
4       DM predicts $\tilde{P}_{n,t}$ for all arm $n\in[N]$ using (15) with $\lambda^{(t)}$.
5       DM computes Whittle Index $W_{n,t}(s)\forall s\in\mathcal{S},n\in[N]$ with $\tilde{P}_{n,t}$ using (7).
@@ 3      All arms starts with state $s_{n,0}$ for all $n\in[N]$
7            DM activates $M$ arms (i.e., action=1) with highest Whittle Indices $W_{n,t}(s_{n,t,h})$.
@@
9             DM observes states $\{s_{n,h+1,t},\forall n\in[N]\}$
10             DM updates counts $C_{t,w_{n}}^{(n)}(s_{n,h+1,t},s_{n,h,t},a_{n,h,t})$ if $(s_{n,h,t},a_{n,h,t})\in\mathcal{Z}_{2}^{(n)}$, otherwise if $(s_{n,h,t},a_{n,h,t})\in\mathcal{Z}_{1}^{(n)}$, DM updates $C_{t,t-1}^{(n)}(s_{n,h+1,t},s_{n,h,t},a_{n,h,t})$
@@ 11      Update $\lambda^{(t+1)}=$ M-th highest Whittle Index
Our sliding window-based online Whittle policy, provided in Algorithm 1, is motivated by the UCWhittle approach proposed in [23]. However, the UCWhittle policy is designed for static settings and does not handle time-varying transition kernels. This motivates the two main technical innovations in our policy. First, we employ a sliding window method that tracks transition kernels of the past $w$ episodes instead of all past episodes (with different sliding windows for different parts of the state space). Second, we utilize two distinct confidence bounds for stationary sets $\mathcal{Z}^{(n)}_{1}$ and non-stationary sets $\mathcal{Z}^{(n)}_{2}$. The window size and the confidence bound for non-stationary sets $\mathcal{Z}^{(n)}_{2}$ are designed as a function of the parameter $\epsilon_{n}$, the upper bound of the divergence rate of transition kernels.
@@ 4.1 Sliding Window and Confidence Bounds
where we define $C^{(n)}_{t,w}(s,a):=\max\bigg{\{}\sum_{s^{\prime}\in\mathcal{S}}C_{t,w}^{(n)}(s^{\prime},a,s),1\bigg{\}}.$
Unlike other upper confidence bound approach [23], we consider two different types of confidence radius; the first one is for the set of states which have unknown but stationary transition probabilities ($\mathcal{Z}_{1}^{(n)}$) and the second one is for states with non-stationary transition probabilities ($\mathcal{Z}_{2}^{(n)}$). We represent these confidence radii as follows
$d_{t,1}^{n}(s,a)=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}},\quad d_{t,2}^{n}(s,a)=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{2}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,w}(s,a)}}+w\epsilon_{n},$ (13)
where $\eta_{1},\eta_{2}>0$, and the sliding window $w\geq 1$ are design parameters. Notice that the term $w\epsilon_{n}$ in the confidence radius $d_{t,2}^{n}(s,a)$ for $\mathcal{Z}^{(n)}_{2}$ measures how far the transition kernels could have drifted over a window of $w$ episodes.
@@
$B_{t}^{(n)}=\bigg{\{}P_{n,t}:\sum_{s^{\prime}\in\mathcal{S}}$ $\bigg{|}P_{n,t}(s^{\prime}|s,a)-\hat{P}_{n,t,w_{i}}(s^{\prime}|s,a)\bigg{|}\leq d_{t,i}^{(n)}(s,a),~{}\forall(s,a)\in\mathcal{Z}^{(n)}_{i},i\in\{1,2\},$
$w_{1}=t,w_{2}=w,P_{n,t}(s^{\prime}|s,a)=0,\forall s^{\prime}\in\mathcal{S}^{(n)}_{0}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}\bigg{\}}.$ (14)
@@ 4.2 Online Whittle Indices
As a result of the maximization procedure (15), the true value function is upper bounded by the value function under the predicted transition kernel provided that the confidence bound (4.1) holds. This upper bound value function will later alllow us to prove regret bounds. Using the predicted transition kernel $\tilde{P}_{n,t}$, we compute $W_{n,t}(s:\tilde{P}_{n,t})$, the Whittle index of state $s\in\mathcal{S}$ for arm $n$ as defined in (7). Finally, we update Lagrange multiplier $\lambda^{t+1}$ as the $M$-th highest Whittle index at time slot $H$ of episode $t$. However, the approach can be computationally intensive due to the kernel maximization problem (15). Some approaches to reduce the complexity are discussed in Appendix D.
@@ 5 Regret Analysis
It is well known in RMAB literature, the problem of obtaining an optimal policy for an RMAB problem is generally intractable [18]. Consequently, using an optimal policy for an RMAB is an impractical benchmark to evaluate the online learning performance. To address this issue, [23] and [34] used the Lagrangian relaxed problem or alternatively the best Whittle index policy to assess the performance of an online learning algorithm. Following [23], we will also employ the Lagrangian relaxed problem for evaluating the online learning performance in this paper.
We determine the regret of the policy $\pi_{t}$ in episode $t$ by subtracting the performance of our policy from the performance of the optimal policy (both under the true unknown transition kernel $P_{n,t}$) for the Lagrange relaxed problem.
$\mathrm{Reg}^{(t)}=\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda^{*})-U(\pi_{n,t},P_{n,t},\lambda^{(t)}),$ (17)
where $\pi^{*}_{n,t}$ is the optimal policy of the Lagrange relaxed problem (4) and $\lambda^{*}$ minimizes Lagrangian objective value $\sum_{n=1}^{N}U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)$. The cumulative dynamic regret in $T$ episodes is given by $\mathrm{Reg}(T)=\sum_{t=1}^{T}\mathrm{Reg}^{(t)}.$ Note that sublinear dynamic regret is usually challenging to establish in online learning literature, since we are comparing to a dynamic optimal policy that knows the entire sequence of transition kernels [36, 37]. We will show that our approach has sublinear dynamic regret, as long as the transition kernels don’t vary too quickly.
To create a regret bound, we first need to establish how good our estimates of the time-varying transition kernel are. To do so, we will bound the probability that the true kernel is outside the high-dimensional ball $B_{t}^{(n)}$ introduced in (4.1). Lemma 1 describes the result in detail.
Lemma 1.
Given $\epsilon_{n},\eta_{1},\eta_{2}\geq 0$, the probability that the true kernel $P_{n,t}$ lies within this high-dimensional Ball $B_{t}^{(n)}$ is greater or equal to $1-\eta$, where $\eta=\eta_{1}+\eta_{2}$, i.e., $\mathrm{Pr}(P_{n,t}\in B_{n,t},\forall n,\forall t)\geq 1-\eta$.
Lemma 1 implies that for every episode $t$, we can provide a confidence region in which true transition kernel will lie with high probability. A detailed proof of Lemma 1 is provided in Appendix A. Next, using this result, Theorem 1 characterizes the dynamic regret for our proposed Algorithm 1.
Theorem 1.
@@
$\mathrm{Reg}(T)\leq O\left(\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|\bigg{)}H\sqrt{T\mathrm{log}T}\right)+\sum_{t=1}^{T}O\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{2}|\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\bigg{)},$ (18)
@@
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{Z}_{2}}g_{t,n}(s,a,w)$ and the function $g_{t,n}(s,a,w)$ is a non-decreasing function given by $g_{t,n}(s,a,w)=\mathbb{E}_{P_{n,t},\pi_{n,t}}[\alpha_{t}^{(n)}(s,a)/\sqrt{C_{t,w}^{(n)}(s,a)}]$, where $\alpha_{t}^{(n)}(s,a)$ is a random variable that denotes the number of visit at $(s,a)$ in episode $t$.
Proof Sketch. We focus on bounding the dynamic regret when the confidence bound holds.
$\mathrm{Reg}^{(t)}$ $=U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)\overset{(a)}{\leq}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)$
$\overset{(b)}{\leq}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{h=1}^{H}\sum_{(s,a)\in\mathcal{S}}\alpha_{t}^{(n)}(s,a)\sum_{s^{\prime}\in\mathcal{S}}\bigg{|}\tilde{P}_{n,t}(s^{\prime}|s,a)-P_{n,t}(s^{\prime}|s,a)\bigg{|}\bigg{]}V_{max},$ (19)
where (a) holds due to (15) and (b) is obtained by using the random variable $\alpha_{t}^{(n)}(s,a)$. Then, we utilize the confidence radii $d_{t,1}^{(n)}(s,a)$, $d_{t,2}^{(n)}(s,a)$, and sum the regret $\mathrm{Reg}^{(t)}$ over episodes $t=1$ to $T$ to get the final regret bound. A detailed proof is provided in Appendix B.
At a high-level, the regret bound consists of two terms - one for parts of the transition kernels that are stationary, and one for parts that are non-stationary. The second term grows with $\epsilon_{n}$ and depends crucially on the choice of window sizes for each arm. Moreover, as the size of the sets $\mathcal{Z}^{(n)}_{1}$ and $\mathcal{Z}^{(n)}_{2}$ increases, the regret increases. If the transition kernels are stationary, i.e., $\epsilon=0$, then our regret becomes $O((\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|)H\sqrt{T\mathrm{log}T}),$ which is better compared to UCWhittle’s [23] regret upper bound $O\left(NH|\mathcal{S}|\sqrt{T\mathrm{log}T}\right)$ since $\mathcal{Z}^{n}_{1}\subseteq\mathcal{S}$. This shows that prior domain knowledge is helpful both for stationary and non-stationary settings.
While Theorem 1 characterizes the regret of our algorithm, the following theorem describes how to optimally choose window sizes for each arm. We are able to show that for slowly varying kernels, our algorithm has sublinear dynamic regret.
Theorem 2.
@@
If there exists a positive probability to visit every $(s,a)\in\mathcal{Z}_{2}$ at least once in every episode $t\in[T]$ for all arms $n\in[N]$ and the divergence of transition dynamics between two consecutive episodes is upper bounded by $\epsilon_{n}=1/T^{k}$, then by using window size $w_{n}=T^{q}$ and $q=\min\{2k/3,1\}$, with probability $1-\eta$, we have $\mathrm{Reg}(T)\leq O(H\sqrt{\mathrm{log}(T)}T^{r})$, where $r=\max\{(3-k)/3,1/2\}$.
Proof Sketch. Given our assumptions, we can upper bound the function $G_{t,n}(w_{n})$ provided in Theorem 1 by $\frac{H}{\sqrt{\mathrm{P}_{min}w}}$, where $\mathrm{P}_{min}$ is the minimum probability to visit any state-action $(s,a)\in\mathcal{Z}_{2}$ at least once at every episode $t\in[T]$. Inserting this bound in Theorem 1 yields Theorem 2. A detailed proof is provided in Appendix C.
6 Numerical Study
Applications $N$ $M$ Our Policy UCWhittle Random WIQL

### v2

@@ 4 Sliding Window-based Online Whittle Policy
To compute the Whittle index, we need to know transition kernels. In practice, transition kernels $P_{n,t}$ are unknown and non-stationary. In this section, we present Algorithm 1, an online approach for RMABs which adapts to unknown and non-stationary transition kernels.
input : State Space $\mathcal{S}$, Action Space $\mathcal{A}$, Reward Function $r_{n}(s,a)$ for all $(s,a)$ and arms $n$
1 DM initializes a Lagrange cost $\lambda^{(1)}$
@@
3  DM predicts variation budget $V_{n}$ for all $n\in[N]$
4   DM decides window size $w_{n}=\lceil(T/V_{n})^{2/3}\rceil$ for all $n\in[N]$
5   Arm $n$ starts with state $s_{n,0}$
6   DM predicts $\tilde{P}_{n,t}$ for all arm $n\in[N]$ using equation 12 with $\lambda^{(t)}$.
7   DM computes Whittle Index $W_{n,t}(s)\forall s\in\mathcal{S},n\in[N]$ with $\tilde{P}_{n,t}$ using equation 7.
@@ 5   Arm $n$ starts with state $s_{n,0}$
9     DM activates $M$ arms (i.e., action=1) with highest Whittle Indices $W_{n,t}(s_{n,h,t})$.
@@
11      DM observes states and updates counts $C_{t,w_{n}}^{(n)}(s_{n,h+1,t},s_{n,h,t},a_{n,h,t})$
12
@@ 13  Update $\lambda^{(t+1)}=$ M-th highest Whittle Index
Our sliding window-based online Whittle policy, provided in Algorithm 1, is motivated by the UCWhittle approach proposed in [30]. However, the UCWhittle policy is designed for static settings and does not handle time-varying transition kernels. This motivates the two main technical innovations in our policy. First, we employ a sliding window method that tracks transition kernels of the past $w_{n}$ episodes instead of all past episodes. The parameter $w_{n}$ is decided based on the total variation budget $V_{n}$. Second, we change the confidence bound provided in [30]. In designing the new confidence bound, we add a prediction horizon $w_{n}V_{n}/T$. We also discuss in Section 4.3 how we estimate the total variation budget $V_{n}$.
@@ 4.1 Sliding Window and Confidence Bounds
where we define $C^{(n)}_{t,w}(s,a):=\max\bigg\{\sum_{s^{\prime}\in\mathcal{S}}C_{t,w}^{(n)}(s^{\prime},a,s),1\bigg\}.$ Using the upper confidence bound approach, we consider the following confidence radius
$d_{t}^{n}(s,a)=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}{C^{(n)}_{t,w}(s,a)}}+\frac{w_{n}V_{n}}{T},$ (10)
where $\eta>0$ is a design parameter. Notice that the term $\frac{w_{n}V_{n}}{T}$ in the confidence radius $d_{t}^{n}(s,a)$ measures how far the transition kernels could have drifted over a window of $w_{n}$ episodes.
@@
$B_{t}^{(n)}=$ $\bigg\{P_{n,t}:\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-\hat{P}_{n,t,w_{i}}(s^{\prime}|s,a)\bigg|\leq d_{t}^{(n)}(s,a),$
$P_{n,t}(s^{\prime}|s,a)=0,\forall s^{\prime}\in\mathcal{S}_{0}(s,a),\sum_{s^{\prime}\in\mathcal{S}}P_{n,t}(s^{\prime}|s,a)=1,\forall(s,a)\in\mathcal{S}\times\mathcal{A}\bigg\}.$ (11)
@@ 4.2 Online Whittle Indices
As a result of the maximization procedure of equation 12, the true value function is upper bounded by the value function under the predicted transition kernel provided that the confidence bound in equation 4.1 holds. This upper bound value function will later allow us to prove regret bounds. Using the predicted transition kernel $\tilde{P}_{n,t}$, we compute $W_{n,t}(s:\tilde{P}_{n,t})$, the Whittle index of state $s\in\mathcal{S}$ for arm $n$ as defined in equation 7. Finally, we update Lagrange multiplier $\lambda^{t+1}$ as the $M$-th highest Whittle index at time slot $H$ of episode $t$. However, the approach can be computationally intensive due to the kernel maximization problem equation 12. Some approaches to reduce the complexity are discussed in Appendix A.5.
4.3 Estimation of Unknown Variation Budget
In the above discussions, the total variation budget $V_{n}$ is assumed to be known. Now, we discuss how to adapt with the unknown variation budget $V_{n}$. We adopt the Bandit over Bandit approach [33, 21] for the estimating the variation budget $V_{n}$. In this estimation approach, we solve another bandit problem to select $V_{n}$ from a finite set of possible budget values based on the history by using EXP3 algorithm [41]. Modified EXP3 algorithm for our problem is provided in Algorithm 2.
Next, we discuss how we can create a set of possible budget values. First, we get the maximum value for the variation budget as $V_{n,max}=2T$. This holds because
$\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\bigg|\leq 2.$ (14)
By using $V_{n,max}$, we can now define the set of qunatized drift values as $\{V_{n,max},V_{n,max}-V_{n,max}/J_{n},V_{n,max}-2V_{n,max}/J_{n},\ldots,V_{n,max}/J_{n}\},$ where $J_{n}$ is the number of quantization levels. This approach of quantizing and approximately estimating drift values is novel within online learning literature. We will show that the number of levels $J_{n}$ affects the dynamic regret (more levels means more accurate tracking of $V_{n}$ but also slower learning in the Bandit-over-Bandit approach).
@@ 5 Regret Analysis
We determine the regret of the policy $\pi_{t}$ in episode $t$ by subtracting the performance of our policy from the performance of the optimal policy (both under the true unknown transition kernel $P_{n,t}$). The cumulative dynamic regret in $T$ episodes is given by
$\mathrm{Reg}(T)$ $=\sum_{t=1}^{T}\bigg(R_{t}\left(\pi^{*}_{t},(P_{n,t})_{n=1}^{N}\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)\bigg)$
$\leq\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)\bigg)$
$=\underbrace{\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)-\sum_{n=1}^{N}U\left(\pi_{n,t},P_{n,t},\lambda\right)\bigg)}_{\mathrm{Term1}}$
$+\underbrace{\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}U\left(\pi_{n,t},P_{n,t},\lambda\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)\bigg)}_{\mathrm{Term2}},$ (15)
where $\pi^{*}_{n,t}$ is the optimal policy of the problem defined in equation 4 associated with transition kernel $P_{n,t}$ and $\pi_{n,t}$ is the optimal policy of the the problem defined in equation 4 associated with transition kernel $\tilde{P}_{n,t}$. The first inequality holds because relaxed Lagrangian upper bounds the main problem.
$\mathrm{Term1}$ is regret on the Lagrangian relaxed problem. To analyze the performance of Whittle index policy, [30] only used the Lagrangian relaxed problem to assess the performance of an online learning algorithm. We consider a stronger version of regret definition by considering $\mathrm{Term2}$ compared to [30]. $\mathrm{Term2}$ is the performance difference between the Lagrangian problem and the original problem with the Whittle index policy derived using the solution of the Lagrangian problem.
First, we analyze $\mathrm{Term1}$. Note that sublinear dynamic regret is usually challenging to establish in online learning literature, since we are comparing to a dynamic optimal policy that knows the entire sequence of transition kernels [42, 43]. We will show that our approach has sublinear dynamic regret, as long as the transition kernels don’t vary too quickly.
To create a regret bound, we first need to establish how good our estimates of the time-varying transition kernel are. To do so, we will bound the probability that the true kernel is outside the high-dimensional ball $B_{t}^{(n)}$ introduced in equation 4.1. Lemma 1 describes the result in detail.
Lemma 1
Given $\eta\geq 0$, the probability that the true kernel $P_{n,t}$ lies within the high-dimensional Ball $B_{t}^{(n)}$ (described by eq. 4.1) is greater than or equal to $1-\eta$, i.e., $\mathrm{Pr}(P_{n,t}\in B_{t}^{(n)},\forall n,\forall t)\geq 1-\eta$.
Lemma 1 implies that for every episode $t$, we can provide a confidence region in which true transition kernel will lie with high probability. A detailed proof of Lemma 1 is provided in Appendix A.1.
Next, using this result, Theorem 1 characterizes the upper bound for $\mathrm{Term1}$.
Theorem 1
@@
$\mathrm{Term1}\leq$ $\sum_{t=1}^{T}O\bigg(\sum_{n=1}^{N}2|\mathcal{S}|G_{t,n}(w_{n})\bigg)+\sum_{n=1}^{N}O\bigg(w_{n}(\tilde{V}_{n}+2T/J_{n})H\bigg)$
$+\sum_{n=1}^{N}O\bigg(\sqrt{J_{n}\mathrm{log}(J_{n})T}\bigg),$ (16)
@@
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}g_{t,n}(s,a,w)$ and the function $g_{t,n}(s,a,w)=\mathbb{E}_{P_{n,t},\pi_{n,t}}[\alpha_{t}^{(n)}(s,a)/\sqrt{C_{t,w}^{(n)}(s,a)}]$ is non-increasing in $w$, where $\alpha_{t}^{(n)}(s,a)$ is a random variable that denotes the number of visit at $(s,a)$ in episode $t$, $J_{n}$ is the number of elements in the set of quantized drift values for arm $n$ and $\tilde{V}_{n}$ is the actual total variation measure, given by
$\tilde{V}_{n}=\sum_{t=1}^{T}\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\bigg|.$ (17)
Proof. See Appendix A.2.
The regret bound for $\mathrm{Term1}$ involves three error components: First term is a transition kernels learning error that decreases with the window size $w_{n}$; Second term is a transition kernels prediction error that increases with both $w_{n}$ and the variation budget $\tilde{V}_{n}$ but decreases with $J_{n}$; Third term is a variation budget learning error that increases with $J_{n}$. The following theorem simplifies this bound.
Theorem 2
@@
If there exists a positive probability to visit every $(s,a)\in\mathcal{S}\times\mathcal{A}$ at least once in any episode $t\in[T]$ for all arms $n\in[N]$ and $w_{n}=\lceil(1/\epsilon_{n})^{2/3}\rceil$, then with probability $1-\eta$, we have
$\mathrm{Term1}\leq\sum_{n=1}^{N}\tilde{O}(T^{2/3}(\tilde{V}_{n}+2T/J_{n})^{1/3})+\tilde{O}(\sqrt{TJ_{n}}).$
Proof. See Appendix A.3.
To develop our final regret bound, we introduce $h(N)$ which is a function of the number of arms NN
$h(N)=\sum_{n=1}^{N}U\left(\pi_{n,t},\tilde{P}_{n,t},\lambda\right)-R_{t}\left(\pi_{t},(\tilde{P}_{n,t})_{n=1}^{N}\right),$ (18)
where $R_{t}(\pi_{t},(P_{n,t})_{n=1}^{N})$ and $U(\pi_{n},P_{n,t},\lambda)$ are defined in equation 2 and equation 4, respectively. The function $h(N)$ represents the gap between the performance of the Lagrangian problem under its optimal solution and the main problem under Whittle index Policy. In $h(N)$, both the optimal solution of Lagrangian problem and the Whittle index Policy are designed and evaluated under same transition kernel $\tilde{P}_{n,t}$. Hence, $h(N)$ does not reflect the learning errors or regret, rather it measures the inherent optimality gap of the Whittle index policy even if we know transition kernels accurately.
Now, we are ready to present the upper bound of the cumulative regret term $\mathrm{Reg}(T)$.
Theorem 3
Under the conditions of Theorem 2, if $J_{n}=O(T^{3/5})$, with probability $1-\eta$, we have
$\mathrm{Reg}(T)$ $\leq\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})+h(N)T,$ (19)
where $\tilde{V}=\max_{n}\tilde{V}_{n}$.
Proof. See Appendix A.4.
Remark 1
According to Theorem 3, the upper bound of $\mathrm{Reg}(T)$ is given by $\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})+h(N)T$, where $\tilde{O}(T^{4/5})$ is the learning error for transition kernels and variation budget. It is proved in [4] that $h(N)=0$ for $N=2$. Other prior works [26, 25, 24, 27] showed that $h(N)\to 0$ as $N\to\infty$ if the RMAB is indexable and has a global attractor point. This suggests that our policy can achieve sub-linear regret $\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})$ for large system size and sub-linear $\tilde{V}$.
6 Simulation Results
Applications $(N,M)$ Our Policy UCWhittle UCWhittle+ Window Random WIQL